# spectrum and local weak convergence of sparse random uniform hypergraphs

### v1

@@ 5. Convergence of $r$-set rooted uniform hypergraphs
This section is devoted to the proof of 3.3. Like the $r=1$ case, the proof of the result is broadly divided in following two propositions.
@@
Proof of 3.3.
The result follows from 5.1 and 5.2. ∎
@@ 5.1. Convergence in expectation
Fix $1\leq r\leq k-1$. We have the following proposition analogous to 4.4.
@@
where $d=\binom{k}{r}-1$ and $\nu_{d,\lambda}$ be the induced measure of $d$-block Galton-Watson tree.
Proof of 5.1 .
The result follows from 5.3 by following the same steps as in the proof of 4.1. We skip the details here. ∎
5.1.1. Proof of 5.3
This subsection is dedicated to prove 5.3. As before, for $1\leq r\leq k-1$, we need to estimate, for $g\in\mathbb{G}^{*}_{w}$,
@@
We have the following lemma to understand the terms $\mu_{n}^{r}(B_{\delta}(g))$ and $\nu_{d,\lambda}(B_{\delta}(g))$. If there is no confusion, for ease of writing, we set $H\in H(n,k,p)$, $\mu_{n}^{r}=\mbox{\bf E}[U_{r}(H)]$ and $G_{r}(H)\equiv G_{r}(H)(S)$. Similar to 4.5 we have the following result.
@@
(29) $\mu_{n}^{r}(B_{\delta}(g))={\bf P}((G_{r}(H),S))_{\lceil\frac{1}{\delta}\rceil}\cong g_{\lceil\frac{1}{\delta}\rceil})$
@@
The result follows by the similar arguments as in the proof of 4.5. We skip the details. ∎
Let $g$ be a rooted weighted graph and $S$ be an $r$-set. Then, we have
@@
5.1.2. An exploration process in an $r$-set rooted hypergrapphs
[…]
@@ 5.1.2. An exploration process in an $r$-set rooted hypergrapphs
$J_{t+1}^{r}=\left\{{\bf v}\in G_{r}(H)\;:\;{\bf v}\subset\bigcup\limits_{e\in R_{t+1}^{r}}e\right\}\cap U_{t}^{r}.$
@@
Since we are considering $k$-uniform hypergraphs, from the definition of $J^{r}_{t}$, it is evident that $|I_{t}^{r}|\leq|R^{r}_{t}|k$. Since $A_{t+1}^{r}=(I_{t+1}^{r}\cup A_{t}^{r})\setminus{\bf v}_{t+1}$ and each $e\in R^{r}_{t+1}$ contains $\bf v^{r}_{t+1}$, we have $|A_{t+1}^{r}|-|A^{r}_{t}|\leq(k-r)|R^{r}_{t+1}|$. Since $|A_{0}^{r}|=r$, using the difference we have
@@
Let $\lambda,k,r,n,p$ be as in Assumption 1 and $\{Y_{t}^{r}\}$ be as defined Equation 31. Then, on an enlarged probability space, there exists a sequence of i.i.d $Z_{1}^{\prime},Z_{2}^{\prime},\ldots,Z_{t_{0}}^{\prime}$ with $Z_{1}^{\prime}\sim\text{Poission}(\lambda)$ such that, for large $n$,
@@
Let $Y_{t}^{r}$ and $Z_{t}^{r}$ be as defined above. Observe that, for $t\leq t_{0}$, we have
@@
Figure 7. An $r$ cardinality set $\bf v^{r}_{t+1}$. Formation of a $k$-uniform hyperedge $e$ (blue shaded) that contains both $\bf v^{r}_{t+1}$ and $j$-elements from $U_{t}^{r}$(the region bounded by the dash-dot rectangle).
@@
which implies that, by 4.7,
@@
For large $n$, $\left[\binom{n-r}{k-r}-\binom{|U_{t}|}{k-r}\right]\approx((n-r)^{k-r}-|U_{t}|^{k-r})\frac{1}{(k-r)!}$. Therefore, by 4.8,
@@
Since $Z_{t}^{r}\sim\text{Binomial}(\binom{n-r}{k-r},p)$, by the maximal coupling and (21), then there exists i.i.d. $Z_{1}^{\prime},\ldots,Z_{t_{0}}^{\prime}$ such that $Z_{t}^{\prime}\sim\text{Poission}(\lambda)$ such that
@@
Let $A_{0}^{\prime}=S$ and $U_{0}^{\prime}=[n]\backslash A_{0}^{\prime}$. For $t\in\mathbb{N}$, we denote
@@
Note that the possible number of edges containing an $r$-set $S$ is $\binom{n-r}{k-r}$. Note that if $|e\backslash U_{t}^{\prime}|=r$ then $e\backslash U_{t}^{\prime}={\bf v}_{t+1}$. Therefore the number of possible edges that satisfies $|e\backslash U_{t}^{\prime}|>r$ is $\binom{n-r}{k-r}-\binom{|U_{t}|}{k-r}$. Thus,
@@
For large $n$, by 4.8, we have
$\left(\binom{n-r}{k-r}-\binom{|U_{t}^{\prime}|}{k-r}\right)$ $\approx(n-r)^{k-r}-|U_{t}^{\prime}|^{k-r}\leq(k-r)(n-r)^{k-r-1}(n-r-|U_{t}^{\prime}|)).$
@@
${\bf P}(|e\backslash U_{t}^{\prime}|>r)$ $\leq\frac{c(n-r-|U_{t}^{\prime}|)}{{(n-r)}}$
$=\frac{c(n-r-\mbox{\bf E}[|U_{t}^{\prime}|])}{n-r}$
@@
Now we ready to give the proof of 5.3.
Proof of 5.3.
@@
$E_{1}^{r}$ $=\{(Z^{\prime}_{1},\ldots,Z^{\prime}_{t_{0}\wedge\tau})\neq(Y_{1}^{r},\ldots,Y^{r}_{t_{0}\wedge\tau})\}.$
$E_{2}^{r}$ $=\{\exists 0\leq s\leq t_{0}\wedge\tau\;:\;|e\backslash U_{t}^{r}|>1\mbox{ for some }e\in R_{t}^{r}\}.$
$E_{3}^{r}$ $=\{\exists 0\leq s\leq t_{0}\wedge\tau\;:\;|e\cap e^{\prime}|>1\mbox{ for some }e\neq e^{\prime}\in R_{t}^{r}\}.$
Using the same arguments as in the proof of 4.4, we have
@@
Thus, combining 5.5, 5.6 and 5.7 and $t_{0}\wedge\tau\leq t_{0}$, we have
@@ 5.2. Almost sure convergence
This section is dedicated to prove 5.2. We follow a similar approach as in the proof of Section 4.2. We prove the result utilizing Chernoff bounds, the Borel-Cantelli lemma, and the Azuma-Hoeffding inequality.
@@
Let $\lambda,k,r,n,p$ be as in Assumption 1, and $\{\mathcal{E}_{j}\}$ be as defined in (26). Suppose $H_{n}\in H(n,k,p)$ and $S_{n}^{r}=U_{r}(H_{n})(A)$ for some fixed Borel set $A\in\mathcal{B}_{m_{r}}$. Then, for $i\leq N$ where $N=\frac{3n\lambda}{2k}$,
@@
Note that $\mbox{\bf E}[|S_{n}^{r}|]\leq 1$. Therefore the Doob’s Martingale Property (1) implies that $\{M_{i}^{r}\}$ is a martingale with respect to the filtration $\{\mathcal{E}_{i}\}$. Observe that
@@
Note that the each term in the right hand side of (37) is bounded above by $\frac{2}{\binom{n}{r}}$. Again, by 4.12, the number hyperedge can be at most $N=\frac{3}{2}\frac{n\lambda}{k}$ almost surely. Thus, if the sigma-algebra $\mathcal{E}_{i}$ exposed the $i$-th hyperedge $e$, then the root vertex belongs to one of the $N$-radius balls centred at some vertex belonging to $e$. Since $e$ contains at most $k$ vertices, the affected $r$-set root is an $r$-set chosen from $c^{\prime}=k\sum\limits_{j=1}^{N}\lambda^{j}$ vertices. Thus, the total number of choices for being an $r$-set root which is affected by revealing the $i$-th hyperedge, is $c=\binom{c^{\prime}}{r}$. Therefore from (37) we have
@@
Proof of 5.2.
5.8 and 2 imply that
@@
where $c_{8}$ is a positive constant (depends on $r,k,\lambda$). By the same arguments as in the proof of Section 4.2, we have
@@
In 4.14, we observed that the condition $\binom{n-1}{k-1}p=\lambda$ can be replaced by $\binom{n-1}{k-1}p\to\lambda$. In the same way, replacing $\binom{n-r}{k-r}p=\lambda$ with $\binom{n-r}{k-r}p\to\lambda$ leads to an identical argument. Hence, 3.3 remains valid under the condition $\binom{n-r}{k-r}p\to\lambda$. Consequently, 3.4 follows from $3.$
@@ 6. Spectral measures associated with hypergraphs
We prove 3.5 in this section. The $r$-set weighted adjacency operator $A_{H}^{w_{r}}:\ell^{2}(\mathcal{V}_{r})\to\ell^{2}(\mathcal{V}_{r})$, defined in (6), coincides with the adjacency matrix of the weighted graph $G_{r}(H)$.
It was shown in [17, Proposition 2.2] that spectral convergence holds for adjacency operators associated with unweighted graphs. To establish 3.5, it therefore suffices to extend that result to the setting of weighted graphs. Our strategy follows the same general approach as in [17, Proposition 2.2], with appropriate modifications to account for the presence of weights. In fact, we refer directly to [17] for the parts of the proof that remain unchanged in the weighted setting.
A probability measure on the space of all rooted graphs is called unimodular when it satisfies the Mass-Transport Principle (see [3, Definition 2.1]). First, we extend this notion for the measures on the space of weighted rooted graphs $\mathbb{G}_{w}^{*}$. A two rooted weighted graphs $(G,w,o,o^{\prime})$ is a weighted graphs with weight $w$ and two roots (distinguished vertices) $o$ and $o^{\prime}$. Two graphs $(G_{1},w_{1},o_{1},o_{1}^{\prime})$ and $(G_{2},w_{2},o_{2},o_{2}^{\prime})$ are called isomorphic if there exists a bijective map $f:V(G_{1})\to V(G_{2})$ such that $(G_{1},w_{1},o_{1})\simeq(G_{2},w_{2},o_{2})$ with respect to $f$ and $f(o_{1}^{\prime})=o_{2}^{\prime}$.
Let $\mathbb{G}^{**}_{w}$ denote the space of all equivalence classes of two rooted locally finite weighted graphs. A Borel probability measure $\mu$ on $\mathbb{G}_{w}^{*}$ is called unimodular if it satisfies the Mass-Transport Principle, that is, for each Borel measurable function $f:\mathbb{G}_{w}^{**}\to[0,\infty)$ the following holds:
@@
The space of all unimodular measures on $\mathbb{G}_{w}^{*}$ is denoted by $\mathcal{P}_{uni}(\mathbb{G}_{w}^{*})$. We have the following reduction, analogous to [3, Proposition 2.2]
@@
Proof of 6.2.
Let $f:\mathbb{G}_{w}^{**}\to[0,\infty)$ be a measurable function such that $f(G,w,u,v)=0$ if $\{u,v\}\notin E(G)$. Observe that, if $u$ is connected with $v$ then $(G(u),w,u,v)=(G(v),w,u,v)$. It follows that
@@
This completes the proof of the result by 6.1. ∎
@@
Since $U_{r}(H_{n})=U(G_{r}(H_{n}))$, where $G_{r}(H)$ is an weighted graph, thus the lemma follows directly from 6.2. ∎
@@
Since $\rho\in\mathcal{P}_{uni}(\mathbb{G}^{*}_{w})$, it satisfy the Mass-Transport Principle
@@
that is, the part of $\mathcal{H}$ supported almost surely on vertices of degree at most $n$. For $\mathbf{f}\in\mathcal{H}_{n}$, the Cauchy–Schwarz inequality leads us to $|(A^{w}_{G}\mathbf{f}_{(G,o)})(v)|^{2}\leq\deg_{G}(v)\|\mathbf{f}_{(G,o)}\|^{2}$. That is, $\rho$-a.s
$\|(A^{w}_{G}P_{n})\mathbf{f}_{(G,o)}\|^{2}\leq\sum\limits_{v\in V(G)}\mathbf{1}_{\{\deg(v)\leq n\}}|((A^{w}_{G}P_{n})\mathbf{f}_{(G,o)})(v)|^{2}\leq n^{2}\|\mathbf{f}_{(G,o)}\|^{2}.$
Thus, $A^{w}_{G}P_{n}$ is bounded operator. Because $A_{G}^{w}P_{n}$ is bounded, it is everywhere defined and extends $A^{w}$ on $\mathcal{H}$ (since $P_{n}f=f$ for $f\in\mathcal{H}_{n}$). Thus, for any $h\in\mathcal{H}_{n}$, we have $A^{w}h=A^{w}P_{n}h,$ which is well-defined and lies in $\mathcal{H}$. Hence $\mathcal{H}_{n}\subset\operatorname{Dom}(\overline{A^{w}}).$
@@
Since both $A^{w}$ and $(A^{w})^{*}$ is affiliated with $\mathcal{M}$, by 6.5, we have $\overline{A^{w}}=(A^{w})^{*}$
@@
Since $A_{n}$ converges to $A$ in the strong resolvent sense, it follows that $f(A_{n})\psi$ converges in norm to $f(A)\psi$ [48, Theorem VIII.20]. Consequently, by continuity,
@@
Thus, for all bounded continuous functions $f$, we have
@@
Observe that if $\bf P$} denotes the probability measure governing the joint distribution of $(G_{n},o)$ and $(G,o)$, then by 6.6, $\bf P$}-a.s. both $A_{n}$ and $A$ are essentially self-adjoint with common core given by the compactly supported functions in $\ell^{2}(V)$. On the other hand, for any compactly supported $f\in\ell^{2}(V)$, one has $A_{n}f=Af$ for all large $n$, where $A_{n}$ and $A$ denote the adjacency operators of $G_{n}$ and $G$, respectively. Therefore $A_{n}$ converges to $A$ in strong resolvent convergence (see [48, Theorem VIII.25 ]).
Then 6.7 implies that a.s. $\mu_{n}^{\mathbf{1}_{o}}$ converges weakly to $\mu^{\mathbf{1}_{o}}$. Taking expectation, we get $\mu_{\rho_{n}}=\mbox{\bf E}\mu_{n}^{\mathbf{1}_{o}}$ converges weakly to $\mu_{\rho}=\mbox{\bf E}\mu^{\mathbf{1}_{o}}$. In fact, The random variables $X_{n}:=\int f\,d\mu^{1_{o}}_{n}$ are all bounded (since $f$ is bounded and $\mu^{1_{o}}_{n}$ are probability measures). Then, one can apply the Bounded Convergence Theorem (BCT):
@@
Finally we give the proof of 3.5.
Proof of 3.5.
@@
That is, $\mu_{\rho_{n}}=\mu_{A}$, the spectral measure of $A$. Since the $r$-set adjacency matrix defined in (6) is an weighted adjacency matrix of the graph $G_{r}(H)$, 3.5 directly follows from 6.8. This completes the proof. ∎

### v2

@@ 5. Convergence of $r$-set rooted uniform hypergraphs
This section is devoted to the proof of Theorem 3.3. Like the $r=1$ case, the proof of the result is broadly divided in following two propositions.
@@
Proof of Theorem 3.3.
The result follows from Proposition 5.1 and Proposition 5.2. ∎
@@ 5.1. Convergence in expectation
Fix $1\leq r\leq k-1$. Note that the following proposition analogous to Proposition 4.4.
@@
where $d=\binom{k}{r}-1$ and $\nu_{d,\lambda}$ denotes the induced measure of $d$-block Galton-Watson tree.
Proof of Proposition 5.1 .
The result follows from Proposition 5.3 by following the same steps as in the proof of Proposition 4.1. We skip the details here. ∎
[…]
This subsection is devoted to the proof of Proposition 5.3. As before, for $1\leq r\leq k-1$, we need to estimate, for $g\in\mathbb{G}^{*}_{w}$,
@@
Observe that the following lemma to understand the terms $\mu_{n}^{r}(B_{\delta}(g))$ and $\nu_{d,\lambda}(B_{\delta}(g))$. If there is no confusion, for ease of writing, we set $H\in H(n,k,p)$, $\mu_{n}^{r}=\mbox{\bf E}[U_{r}(H)]$ and $G_{r}(H)\equiv G_{r}(H)(S)$. Similar to Lemma 4.5 we have the following result.
@@
(30) $\mu_{n}^{r}(B_{\delta}(g))={\bf P}\left((G_{r}(H),S)_{\lceil\frac{1}{\delta}\rceil}\cong g_{\lceil\frac{1}{\delta}\rceil}\right)$
@@
The result follows by the similar arguments as in the proof of Lemma 4.5. We skip the details. ∎
Let $g$ be a rooted weighted graph and $S$ be an $r$-set. Then,
@@
5.1.2. An exploration process in an $r$-set rooted hypergraphs
We use the breadth-first search exploration procedure as in Section 4.1.2. Let $\mathcal{S}\subset\mathbb{N}_{d}^{\infty}$, where $d=\binom{k}{r}-1$. Define a bijection $\phi:\mathcal{S}\to V(G_{r}(H))$, where $V(G_{r}(H))$ denotes the vertex set of the connected component of $G_{r}(H)$ containing the $r$-set root $S$. Given a vertex ${\bf j}\in V(G_{r}(H))$, the pre-image $\phi^{-1}({\bf j})$ being defined means that $\bf j$} has been explored.
@@ 5.1.2. An exploration process in an $r$-set rooted hypergraphs
$J_{t+1}^{r}=\left\{{\bf v}\in V(G_{r}(H))\;:\;{\bf v}\subset\bigcup\limits_{e\in R_{t+1}^{r}}e\right\}\cap U_{t}^{r}.$
@@
Since we are considering $k$-uniform hypergraphs, from the definition of $J^{r}_{t}$, it is evident that $|I_{t}^{r}|\leq|R^{r}_{t}|k$. Since $A_{t+1}^{r}=(I_{t+1}^{r}\cup A_{t}^{r})\setminus{\bf v}_{t+1}$ and each $e\in R^{r}_{t+1}$ contains $\bf v^{r}_{t+1}$, it follows that $|A_{t+1}^{r}|-|A^{r}_{t}|\leq(k-r)|R^{r}_{t+1}|$. Since $|A_{0}^{r}|=r$, using the difference we have
@@
Let $\lambda,k,r,n,p$ be as in Assumption 1 and $\{Y_{t}^{r}\}$ be as defined Equation 32. Then, on an enlarged probability space, there exists a sequence of i.i.d $Z_{1}^{\prime},Z_{2}^{\prime},\ldots,Z_{t_{0}}^{\prime}$ with $Z_{1}^{\prime}\sim\text{Poisson}(\lambda)$ such that, for large $n$,
@@
Let $Y_{t}^{r}$ and $Z_{t}^{r}$ be as defined above. Observe that, for $t\leq t_{0}$, the following holds:
@@
Figure 7. An $r$-set $\bf v^{r}_{t+1}$. Formation of a $k$-uniform hyperedge $e$ (blue shaded) that contains both $\bf v^{r}_{t+1}$ and $j$-elements from $U_{t}^{r}$(the region bounded by the dash-dot rectangle).
@@
which implies that, by Lemma 4.10,
@@
For large $n$, $\left[\binom{n-r}{k-r}-\binom{|U_{t}|}{k-r}\right]\approx((n-r)^{k-r}-|U_{t}|^{k-r})\frac{1}{(k-r)!}$. Therefore, by Lemma 4.11,
@@
Since $Z_{t}^{r}\sim\text{Binomial}(\binom{n-r}{k-r},p)$, by the maximal coupling and (22), then there exists i.i.d. $Z_{1}^{\prime},\ldots,Z_{t_{0}}^{\prime}$ such that $Z_{t}^{\prime}\sim\text{Poisson}(\lambda)$ such that
@@
Let $A_{0}^{r}=S$ and $U_{0}^{r}=[n]\backslash A_{0}^{\prime}$. For $t\in\mathbb{N}$, we denote
@@
Note that the possible number of edges containing an $r$-set $S$ is $\binom{n-r}{k-r}$. Note that if $|e\backslash U_{t}^{r}|=r$ then $e\backslash U_{t}^{r}={\bf v}_{t+1}$. Therefore the number of possible edges that satisfies $|e\backslash U_{t}^{r}|>r$ is $\binom{n-r}{k-r}-\binom{|U_{t}^{r}|}{k-r}$. Thus,
@@
For large $n$, by Lemma 4.11, note that
$\left(\binom{n-r}{k-r}-\binom{|U_{t}^{r}|}{k-r}\right)$ $\approx(n-r)^{k-r}-|U_{t}^{r}|^{k-r}\leq(k-r)(n-r)^{k-r-1}(n-r-|U_{t}^{r}|)).$
@@
${\bf P}(|e\backslash U_{t}^{r}|>r)$ $\leq\frac{c(n-r-|U_{t}^{r}|)}{{(n-r)}}$
$=\frac{c(n-r-\mbox{\bf E}[|U_{t}^{r}|])}{n-r}$
@@
Now we are ready to give the proof of Proposition 5.3.
Proof of Proposition 5.3.
@@
$E_{1}^{r}$ $=\{(Z^{\prime}_{1},\ldots,Z^{\prime}_{t_{0}\wedge t})\neq(Y_{1}^{r},\ldots,Y^{r}_{t_{0}\wedge t})\}.$
$E_{2}^{r}$ $=\{\exists 0\leq s\leq t_{0}\wedge t\;:\;|e\backslash U_{t}^{r}|>1\mbox{ for some }e\in R_{t}^{r}\}.$
$E_{3}^{r}$ $=\{\exists 0\leq s\leq t_{0}\wedge t\;:\;|e\cap e^{\prime}|>1\mbox{ for some }e\neq e^{\prime}\in R_{t}^{r}\}.$
Using the same arguments as in the proof of Proposition 4.4, we have
@@
Thus, combining Lemma 5.5, Lemma 5.6 and Lemma 5.7 and $t_{0}\wedge t\leq t_{0}$,
@@ 5.2. Almost sure convergence
This section is dedicated to proving Proposition 5.2. We follow a similar approach as in the proof of Proposition 4.2. We prove the result utilizing Chernoff bounds, the Borel-Cantelli lemma, and the Azuma-Hoeffding inequality.
@@
Let $\lambda,k,r,n,p$ be as in Assumption 1, and $\{\mathcal{E}_{j}\}$ be as defined in (27). Let $H_{n}\in H(n,k,p)$ and define $S_{n}^{r}=U_{r}(H_{n})(A)$ for some fixed Borel set $A\in\mathcal{B}_{m_{r}}$. Then, for $i\leq N$ where $N=\frac{3n\lambda}{2k}$,
@@
Note that $\mbox{\bf E}[|S_{n}^{r}|]\leq 1$. Therefore the Doob’s Martingale Property (Fact 1) implies that $\{M_{i}^{r}\}$ is a martingale with respect to the filtration $\{\mathcal{E}_{i}\}$. Observe that
@@
Note that each term in the right hand side of (38) is bounded above by $\frac{2}{\binom{n}{r}}$. Again, by Lemma 4.15, the number hyperedge can be at most $N=\frac{3}{2}\frac{n\lambda}{k}$ almost surely. Thus, if the sigma-algebra $\mathcal{E}_{i}$ exposed the $i$-th hyperedge $e$, then the root vertex belongs to one of the $N$-radius balls centred at some vertex belonging to $e$. Since $e$ contains at most $k$ vertices, the affected $r$-set root is an $r$-set chosen from $c^{\prime}=k\sum\limits_{j=1}^{N}\lambda^{j}$ vertices. Thus, the total number of choices for being an $r$-set root which is affected by revealing the $i$-th hyperedge, is $c=\binom{c^{\prime}}{r}$. Therefore from (38) it follows that
@@
Proof of Proposition 5.2.
Lemma 5.8 and Fact 2 imply that
@@
where $c_{8}$ is a positive constant (depends on $r,k,\lambda$). By the same arguments as in the proof of Proposition 4.2, we have
@@
In Remark 4.17, we observed that the condition $\binom{n-1}{k-1}p=\lambda$ can be replaced by $\binom{n-1}{k-1}p\to\lambda$. In the same way, replacing $\binom{n-r}{k-r}p=\lambda$ with $\binom{n-r}{k-r}p\to\lambda$ leads to an identical argument. Hence, Theorem 3.3 remains valid under the condition $\binom{n-r}{k-r}p\to\lambda$. Consequently, Corollary 3.4 follows from Theorem 3.3.
@@ 6. Spectral measures associated with hypergraphs
We prove Theorem 3.5 in this section. The $r$-set weighted adjacency operator $A_{H}^{w_{r}}:\ell^{2}(\mathcal{V}_{r})\to\ell^{2}(\mathcal{V}_{r})$, defined in (6), coincides with the adjacency matrix of the weighted graph $G_{r}(H)$.
It was shown in [17, Proposition 2.2] that spectral convergence holds for adjacency operators associated with unweighted graphs. That proof rely only on symmetry and boundedness assumptions, which remain true in the weighted case. To establish Theorem 3.5, it therefore suffices to extend that result to the setting of weighted graphs. For this extension we mainly use the fact that in the weighted case, weights only affect edge-supported functions multiplicatively, not structurally. Our strategy follows the same general approach as in [17, Proposition 2.2], with appropriate modifications to account for the presence of weights. In fact, we refer directly to [17] for the parts of the proof that remain unchanged in the weighted setting.
A probability measure on the space of all rooted graphs is called unimodular when it satisfies the Mass-Transport Principle (see [3, Definition 2.1]). First, we extend this notion for the measures on the space of weighted rooted graphs $\mathbb{G}_{w}^{*}$. A two-rooted weighted graphs $(G,w,o,o^{\prime})$ is a weighted graphs with weight $w$ and two roots (distinguished vertices) $o$ and $o^{\prime}$. Two graphs $(G_{1},w_{1},o_{1},o_{1}^{\prime})$ and $(G_{2},w_{2},o_{2},o_{2}^{\prime})$ are called isomorphic if there exists a bijective map $f:V(G_{1})\to V(G_{2})$ such that $(G_{1},w_{1},o_{1})\simeq(G_{2},w_{2},o_{2})$ with respect to $f$ and $f(o_{1}^{\prime})=o_{2}^{\prime}$.
Let $\mathbb{G}^{**}_{w}$ denote the space of all equivalence classes of two-rooted locally finite weighted graphs. A Borel probability measure $\mu$ on $\mathbb{G}_{w}^{*}$ is called unimodular if it satisfies the Mass-Transport Principle, that is, for each Borel measurable function $f:\mathbb{G}_{w}^{**}\to[0,\infty)$ the following holds:
@@
The space of all unimodular measures on $\mathbb{G}_{w}^{*}$ is denoted by $\mathcal{P}_{uni}(\mathbb{G}_{w}^{*})$. Note that the following reduction, analogous to [3, Proposition 2.2]
@@
Proof of Lemma 6.2.
Let $f:\mathbb{G}_{w}^{**}\to[0,\infty)$ be a measurable function such that $f(G,w,u,v)=0$ if $\{u,v\}\notin E(G)$. Observe that, if ${u,v}\in E(G)$ then $(G(u),w,u,v)=(G(v),w,u,v)$. It follows that
@@
This completes the proof of the result by Lemma 6.1. ∎
@@
Since $U_{r}(H_{n})=U(G_{r}(H_{n}))$, where $G_{r}(H)$ is an weighted graph, thus the lemma follows directly from Lemma 6.2. ∎
@@
Since $\rho\in\mathcal{P}_{uni}(\mathbb{G}^{*}_{w})$, it satisfies the Mass-Transport Principle
@@
For $\mathbf{f}\in\mathcal{H}_{n}$ and $v\in V(G)$, we have
$(A^{w}_{G}\mathbf{f})(v)=\sum_{u\sim v}w(u,v)\,\mathbf{f}(u).$
By the Cauchy–Schwarz inequality,
$\big|(A^{w}_{G}\mathbf{f})(v)\big|^{2}\;\leq\;\left(\sum_{u\sim v}|w(u,v)|^{2}\right)\left(\sum_{u\sim v}|\mathbf{f}(u)|^{2}\right)\;\leq\;\left(\sum_{u\sim v}|w(u,v)|^{2}\right)\|\mathbf{f}\|^{2}.$
Thus, $\rho$-a.s.
$\|(A^{w}_{G}P_{n})\mathbf{f}\|^{2}$ $=\sum_{v\in V(G)}\mathbf{1}_{\{\deg(v)\leq n\}}\,\big|(A^{w}_{G}\mathbf{f})(v)\big|^{2}$
$\leq\sum_{v\in V(G)}\mathbf{1}_{\{\deg(v)\leq n\}}\left(\sum_{u\sim v}|w(u,v)|^{2}\right)\|\mathbf{f}\|^{2}.$
If the weights are uniformly bounded, i.e.
$M:=\sup_{u,v}|w(u,v)|<\infty,$
then
$\sum_{u\sim v}|w(u,v)|^{2}\leq M^{2}\deg(v)\leq M^{2}n,$
and therefore
$\|(A^{w}_{G}P_{n})\mathbf{f}\|^{2}\;\leq\;(M^{2}n^{2})\,\|\mathbf{f}\|^{2}.$
Hence, $A^{w}_{G}P_{n}$ is a bounded operator with operator norm at most $Mn$.
Because $A_{G}^{w}P_{n}$ is bounded, it is everywhere defined and extends $A^{w}$ on $\mathcal{H}$ (since $P_{n}f=f$ for $f\in\mathcal{H}_{n}$). Thus, for any $h\in\mathcal{H}_{n}$, note that $A^{w}h=A^{w}P_{n}h,$ which is well-defined and lies in $\mathcal{H}$. Hence $\mathcal{H}_{n}\subset\operatorname{Dom}(\overline{A^{w}}).$
@@
Since both $A^{w}$ and $(A^{w})^{*}$ is affiliated with $\mathcal{M}$, by Lemma 6.5, we have $\overline{A^{w}}=(A^{w})^{*}$
@@
Since $A_{n}$ converges to $A$ in the strong resolvent sense, it follows that $f(A_{n})\psi$ converges in norm to $f(A)\psi$ [49, Theorem VIII.20]. Consequently, by continuity,
@@
Thus, for all bounded continuous functions $f$, observe that
@@
Observe that if $\bf P$} denotes the probability measure governing the joint distribution of $(G_{n},o)$ and $(G,o)$, then by Lemma 6.6, $\bf P$}-a.s. both $A_{n}$ and $A$ are essentially self-adjoint with common core given by the compactly supported functions in $\ell^{2}(V)$. On the other hand, for any compactly supported $f\in\ell^{2}(V)$, one has $A_{n}f=Af$ for all large $n$, where $A_{n}$ and $A$ denote the adjacency operators of $G_{n}$ and $G$, respectively. Therefore $A_{n}$ converges to $A$ in strong resolvent convergence (see [49, Theorem VIII.25 ]).
Then Lemma 6.7 implies that a.s. $\mu_{n}^{\mathbf{1}_{o}}$ converges weakly to $\mu^{\mathbf{1}_{o}}$.
Taking expectation, we get $\mu_{\rho_{n}}=\mbox{\bf E}\mu_{n}^{\mathbf{1}_{o}}$ converges weakly to $\mu_{\rho}=\mbox{\bf E}\mu^{\mathbf{1}_{o}}$. In fact, The random variables $X_{n}:=\int f\,d\mu^{1_{o}}_{n}$ are all bounded (since $f$ is bounded and $\mu^{1_{o}}_{n}$ are probability measures). Then, one can apply the Bounded Convergence Theorem (BCT):
@@
Finally we give the proof of Theorem 3.5.
Proof of Theorem 3.5.
@@
That is, $\mu_{\rho_{n}}=\mu_{A}$, the spectral measure of $A$. Since the $r$-set adjacency matrix defined in (6) is an weighted adjacency matrix of the graph $G_{r}(H)$, Theorem 3.5 directly follows from Lemma 6.8. This completes the proof. ∎